R is the midpoint of PS.Q is the midpoint of PR. PR=3. what is the length of QS
1
Expert's answer
2020-11-02T20:00:07-0500

Since "Q" is the midpoint of "PR" , "PR=2QR", and thus "QR=\frac{PR}{2}=\frac{3}{2}=1.5". Taking into account that "R" is the midpoint of "PS" we conclude that "RS=PR=3". Therefore, "QS=QR+RS=1.5+3=4.5."